A fire is raging in an apartment house. You see kids in a window in an upstairs window. The window is 24feet off the ground. You have a ladders, but it only extends to only 30 feet. How far from the building must you place the ladder to reach the kids?

Please solve and answer. Thank you

Answer by edjones(8007)

a^2+b^2=c^2
24^2+b^2=30^2
b^2=30^2-24^2
b^2=324
b=18' dist from building to place ladder.
.
Ed